OpenStudy (anonymous):

A culture started with 6,000 bacteria After 2 hours, it grew to 7,800 bacteria. Predict how many bacteria will be present after 15 hours. Round your answer to the nearest whole number. I think the answer is 19500.

OpenStudy (anonymous):

7800 - 6000=18000 1800/7.5 =13500 13500+6000=19500

OpenStudy (anonymous):

7800 - 6000=18000 1800/7.5 =13500 13500+6000=19500